OpenStudy (anonymous):

Help with geometric sequence please! Find the nth term for the following geometric sequence: t4 = 8x^3 and t9 = 5x^20

OpenStudy (anonymous):

\[t9=8x^{3} and t9 = 5x^{20}\]

OpenStudy (anonymous):

this is what I have: t4 = 8x^3 8x^3 = ar^3 a = 8x^3 / r^3 5x^20 = ar^19 5x^20 = (8x^3)(r^19) / r^3 5x^20 = (8x^3)(r^16) 5x^20 / 8x^3 = (8x^3)(r^3) / 8x^3 0.625x^17 = r^16 Now I'm stuck.

OpenStudy (john_es):

In the problem, is t_{9} or t_{19}?

OpenStudy (anonymous):

t9 = term 9. the answer is supposed to e tn = (2x)^n-1 but I don't know how to get it

OpenStudy (john_es):

Then, I thin it should be this way, \[t_n=t_1r^{n-1}\]\[t_4=t_1r^3=8x^3\\ t_9=t_1r^8=5x^{20}\] Divide both expressions, to find r, \[\frac{t_1r^8}{t_1r^3}=\frac{5x^{20}}{8x^3}\] Then use any of the before equations t_4 or t_9 to find t_1. At last, you should do t_n=t_1*r^{n-1}. Try it.

OpenStudy (john_es):

However, the solution you give is not the correct solution for the problem you posted. You can check it. With the solution you give, \[t4=8\cdot x^3\\ t9=256\cdot x^8\] So or the problem is wrong or the solution is wrong.